OpenStudy (anonymous):

Please helppppppp triple sad face!!!! Elimination method. 5x + 2y = 55 2x - y = 19

OpenStudy (anonymous):

you want to get x or y?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

this one is tricky very tricky

OpenStudy (anonymous):

okay so you have to work on one equation and have only one variable go to the second equation and move y to the other place it will be y=2x-19 so now you can eliminate y in the first equation with 2x-19 it will be 5x+2(2x-19)=55 simplify .. 5x + 4x-38+55 >> 9x=93 then x equals 31/3

OpenStudy (anonymous):

hang on lemme try this by myself too

OpenStudy (anonymous):

wait you lost me

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

the answer is x + y= 12 i need to know why

OpenStudy (anonymous):

sorry i lost connection well in case you havent known it yet we got x= 31/3 then eliminate in the second equation ... 2(31/3) - y = 19 simplify ... 62/3 - y =19 then switch places of y and 19 with opp signs to get y like that 62/2 - 19 = y then y equal 5/3

OpenStudy (anonymous):

now you want to know x + y x = 31/3 y=5/3 so substitute 31/3 + 5/3 equals 12

OpenStudy (anonymous):

i hope its clear now :)